OpenStudy (anonymous):

Find the remainder when 3^1993 + 11 is divided by 9.

OpenStudy (anonymous):

\[3^{1993}+11 \] is divided by 9.

OpenStudy (blockcolder):

We know that \[a \equiv b \pmod{m}\] means that the remainder when a is divided by m is b. We have \[3^{1993} \equiv 3 \pmod{9}\\ 11 \equiv 2 \pmod{9}\] Therefore, \[3^{1993}+11 \equiv 5 \pmod{9}\] and the remainder is 5.

OpenStudy (blockcolder):

That first one should be 0, I mean. =))

OpenStudy (anonymous):

Can it be solved without using mod? We didn't do it.

OpenStudy (blockcolder):

Well, we can do it like this: By the division algorithm, we have \[3^{1993}+11=9k+r \qquad 0\leq r < 9\] where r is the remainder. We can find r by doing the ff: \[3^{1993}+11=9(3^{1991})+9+2=9(3^{1991}+1)+2\] Comparing this with the above, it can be seen that the remainder is 2.

OpenStudy (anonymous):

But didn't you calculate 5 with the previous method?

OpenStudy (blockcolder):

The first line was wrong. It was supposed to be \[3^{1993}\equiv 0 \pmod{9}\] and adding it to the second line there will give us: \[3^{1993}+11 \equiv 2 \pmod{9}\]

OpenStudy (anonymous):

Okay. Thank you so much!

OpenStudy (blockcolder):

You're welcome. :D